# Gibbs sampling, exact sampling, and distribution function evaluation for the generalized inverse Gaussian distribution

### v2

@@ Abstract
The generalized inverse Gaussian, denoted $\mathrm{GIG}(p,a,b)$, is a flexible family of distributions that includes the gamma, inverse gamma, and inverse Gaussian distributions as special cases. In addition to its applications in statistical modeling and its theoretical interest, the GIG often arises in computational statistics, especially in Markov chain Monte Carlo (MCMC) algorithms for posterior inference. This article introduces two mixture representations for the GIG: one that expresses the distribution as a continuous mixture of inverse Gaussians and another that reveals a recursive relationship between GIGs with different values of $p$. The former representation forms the basis for a data augmentation scheme that leads to a geometrically ergodic Gibbs sampler for the GIG. This simple Gibbs sampler, which alternates between gamma and inverse Gaussian conditional distributions, can be incorporated within an encompassing MCMC algorithm when simulation from a GIG is required. The latter representation leads to algorithms for exact, rejection-free sampling as well as CDF evaluation for the GIG with half-integer $p.$ We highlight computational examples from the literature where these new algorithms could be applied.
@@ 1 Introduction
where $K_{p}(\cdot)$ is the modified Bessel function of the second kind [1]. The moments of the $\mathrm{GIG}(p,a,b)$ involve Bessel functions but are available in closed form (see, for example, [22]).
The generalized inverse Gaussian was first proposed by Halphen in the 1940s for analyzing hydrological data [15, 30]. It regained popularity in the mid to late 1970s, when a series of articles studied its properties in detail (see, for example, [5], [8], and [14]). For an account of known properties and characterizations of the $\mathrm{GIG}(p,a,b)$, we refer the reader to the monograph [22] and the review article [24].
The GIG$(p,a,b)$ often arises in computational statistics, especially in Markov chain Monte Carlo (MCMC) algorithms for posterior inference. For example, it emerges as a full conditional distribution for scale parameters in normal models. Sections 2 and 3 describe or reference a variety of applications where it is necessary to simulate random draws from the $\mathrm{GIG}(p,a,b)$ or evaluate its cumulative distribution function (CDF).
The generalized inverse Gaussian is challenging to simulate from or evaluate the CDF of. Several articles have been written on generating random draws from the $\mathrm{GIG}(p,a,b)$, including [11], [25], [19], [12], and [32]. Recently, [31] derived expressions for the CDF of the generalized inverse Gaussian for half-integer $p$.
This article introduces two novel mixture representations of the generalized inverse Gaussian. One expresses the $\mathrm{GIG}(p,a,b)$ as a continuous mixture of inverse Gaussians, while the other reveals a relationship between GIGs with different values of $p$. The former representation forms the basis for a data augmentation scheme that leads to a geometrically ergodic Gibbs sampler for the GIG. This simple Gibbs sampler, which alternates between gamma and inverse Gaussian conditional distributions, can be incorporated within an encompassing MCMC algorithm when simulation from a GIG is required. The latter representation leads to the first algorithm in the literature for exact, rejection-free sampling from the GIG with half-integer $p$ as well as an algorithm for CDF evaluation in that same setting. The algorithms are included in the article and can be easily implemented in standard programming languages. All proofs appear in the Supplementary Material.
@@ 1.1 Notation
We use the notation $\text{InvGauss}(\mu,\lambda)$ for the inverse Gaussian distribution with mean parameter $\mu$ and shape parameter $\lambda$, $\text{Gamma}(\alpha,\beta)$ for the gamma distribution with shape parameter $\alpha$ and rate parameter $\beta$, $\text{Exp}(\theta)$ for the exponential distribution with rate parameter $\theta$, $\text{Normal}(\mu,\sigma^{2})$ for the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and $\text{Unif}(0,1)$ for the uniform distribution between zero and one. We denote the PDF of the $\mathrm{GIG}(p,a,b)$ evaluated at $x$ by $\mathrm{GIG}(x\mid p,a,b)$. We use analogous notation for the PDFs of other distributions.
@@ 2 Data-augmented Gibbs sampler
In this section, we introduce a representation of the GIG$(p,a,b)$ as a continuous mixture of inverse Gaussian distributions. The representation is useful for deriving a data-augmented Gibbs sampler for which the stationary distribution of one of the marginals is GIG$(p,a,b)$ (Algorithm 1). We then reference examples from the literature where the GIG appears as a full-conditional distribution in Gibbs sampling algorithms. Example 1 shows how Algorithm 1 can be nested within an encompassing MCMC algorithm when simulation from a GIG is required. Finally, we provide theoretical support for the Gibbs sampler described in Algorithm 1. In particular, Proposition 3 establishes that it is geometrically ergodic.
@@
where $f_{X}(x)$ is the PDF of the $\mathrm{GIG}(p,a,b)$, $f_{Y}(y)$ is a PDF with support on $\mathbb{R}_{+}$, and $f_{X\mid Y}(x\mid y)$ is an inverse Gaussian PDF. The parameters of the distributions depend on whether $p<-1/2$ or $p>-1/2$. If $p<-1/2$, then
$f_{X\mid Y}(x\mid y)$ $=\mathrm{InvGauss}(x\mid\sqrt{b/(a+2y)},b)$

### v3

# Properties of the generalized inverse Gaussian with applications to Monte Carlo simulation and distribution function evaluation

@@ Abstract
The generalized inverse Gaussian, denoted $\mathrm{GIG}(p,a,b)$, is a flexible family of distributions that includes the gamma, inverse gamma, and inverse Gaussian distributions as special cases. In addition to its applications in statistical modeling and its theoretical interest, the GIG often arises in computational statistics, especially in Markov chain Monte Carlo (MCMC) algorithms for posterior inference. This article introduces two mixture representations for the GIG: one that expresses the distribution as a continuous mixture of inverse Gaussians and another that reveals a recursive relationship between GIGs with different values of $p$. The former representation forms the basis for a data augmentation scheme that leads to a geometrically ergodic Gibbs sampler for the GIG. This simple Gibbs sampler, which alternates between gamma and inverse Gaussian conditional distributions, can be incorporated within an encompassing MCMC algorithm when simulation from a GIG is required. The latter representation leads to algorithms for exact, rejection-free sampling as well as CDF evaluation for the GIG with half-integer $p.$
@@ 1 Introduction
where $K_{p}(\cdot)$ is the modified Bessel function of the second kind (Abramowitz and Stegun, 1968). The moments of the GIG involve Bessel functions but are available in closed form (see, for example, Jørgensen (1982)).
The generalized inverse Gaussian was first proposed by Halphen in the 1940s for analyzing hydrological data (Halphen, 1941; Perreault et al., 1999). It regained popularity in the mid to late 1970s, when a series of articles studied its properties in detail (see, for example, Barndorff-Nielsen and Halgreen (1977), Blæsild (1978), and Halgreen (1979)). For an account of known properties and characterizations of the $\mathrm{GIG}(p,a,b)$, we refer the reader to the monograph Jørgensen (1982) and the review article Koudou and Ley (2014).
The GIG often arises in computational statistics, especially in Markov chain Monte Carlo (MCMC) algorithms for posterior inference. For example, it emerges as a full conditional distribution for scale parameters in normal models. Sections 2 and 3 describe or reference a variety of applications where it is necessary to simulate random draws from the GIG or evaluate its cumulative distribution function (CDF).
This article introduces two mixture representations of the generalized inverse Gaussian. One expresses the $\mathrm{GIG}(p,a,b)$ as a continuous mixture of inverse Gaussians, while the other reveals a relationship between GIGs. The former representation forms the basis for a data augmentation scheme that leads to a geometrically ergodic Gibbs sampler for the GIG. This Gibbs sampler, which alternates between gamma and inverse Gaussian conditional distributions, can be incorporated within an encompassing MCMC algorithm when simulation from a GIG is required. The latter representation leads to the first algorithm in the literature for exact, rejection-free sampling from the GIG with half-integer $p$ as well as an algorithm for CDF evaluation in that same setting.
Several articles have been written on the topic of simulating from the GIG, including Dagpunar (1989), Hörmann and Leydold (2014), Devroye (2014), and Zhang and Reiter (2022). Recently, Willmot and Woo (2022) derived expressions for the CDF of the generalized inverse Gaussian for half-integer $p$. In simulation experiments, we found that our implementations of the data-augmented Gibbs sampler and the exact simulation method can be more efficient than prominent R packages for simulating from the GIG, but this is not always the case. In our view, the methods proposed in this article complement rather than supersede the existing methods.
The proofs of all propositions can be found in the supplementary material. In addition, the supplementary material includes pseudocode for the algorithms, a section discussing an alternative Gibbs sampler that can be derived from the results of Zhang and Reiter (2022), and a section describing how to adapt the propositions to an alternative parametrization of the GIG. The supplementary material also contains a discussion of the simulation experiments [^1] Code available at https://github.com/michaeljauch/gig..
@@ 1.1 Notation
We use the notation $\text{InvGauss}(\mu,\lambda)$ for the inverse Gaussian distribution with mean parameter $\mu$ and shape parameter $\lambda$. Similarly, $\text{Gamma}(\alpha,\beta)$ is the gamma distribution with shape parameter $\alpha$ and rate parameter $\beta$ and $\text{Exp}(\theta)$ is the exponential distribution with rate parameter $\theta$.
@@ 2 Data-augmented Gibbs sampler
In this section, we introduce a representation of the GIG$(p,a,b)$ as a continuous mixture of inverse Gaussian distributions. The representation is useful for deriving a data-augmented Gibbs sampler for the GIG. We then reference examples from the literature where the GIG appears as a full-conditional distribution in Gibbs sampling algorithms. An example shows how the Gibbs sampler can be nested within an encompassing MCMC algorithm when simulation from a GIG is required. Finally, we provide theoretical support for the Gibbs sampler: Proposition 3 establishes that it is geometrically ergodic.
@@
where $f_{X}(x)$ is the PDF of the $\mathrm{GIG}(p,a,b)$, $f_{Y}(y)$ is a PDF with support on $\mathbb{R}_{+}$, and $f_{X\mid Y}(x\mid y)$ is an inverse Gaussian PDF.
The parameters of the distributions are different for the cases $p<-1/2$ and $p>-1/2$. If $p<-1/2$, then
$f_{X\mid Y}(x)$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,=f_{\mathrm{IG}}(x\,;\,\sqrt{b/(a+2y)},b)}$
$\mathcal{Z}_{0}$ $=\frac{b^{(p+1)/2}\,\sqrt{2}\,\Gamma(-p-1/2)\,K_{p}(\sqrt{ab})}{a^{p/2}\,\sqrt{\pi}},$
where $f_{\mathrm{IG}}(x\,;\,\sqrt{b/(a+2y)},b)$ is the PDF of the InvGauss$(\sqrt{b/(a+2y)},b)$. If $p>-1/2$, then
$f_{X\mid Y}(x)$ $\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=f_{\mathrm{IG}}(x\,;\,\sqrt{(b+2y)/a},b+2y)}$
To see why we have two separate cases in Proposition 1, notice that the PDF of the GIG can be written as
$f_{X}(x)\propto x^{p+1/2}\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{\text{IG}}(x\,;\,\sqrt{b/a},b)}.$
We can substitute two different integral representations for $x^{p+1/2}:$ one that is valid for $p>-1/2$ and another one that is valid for $p<-1/2.$ If $p>-1/2$, we use the representation
$x^{p+1/2}\underset{(p>-1/2)}{=}\frac{\int_{0}^{\infty}y^{p-1/2}\,\exp(-y/x)\,\mathrm{d}y}{\Gamma(p+1/2)}.$
If $p<-1/2$, we use the representation
$x^{p+1/2}\underset{(p<-1/2)}{=}\frac{\int_{0}^{\infty}y^{-(p+3/2)}\,\exp\left(-xy\right)\,\mathrm{d}y}{\Gamma(-(p+1/2))}.$
The case $p=-1/2$ is not included in Proposition 1, but the GIG$(-1/2,a,b)$ is equivalent to $\mathrm{InvGauss}(\sqrt{b/a},b)$ (Jørgensen, 1982). The conditional distribution $X\mid Y=y$ can be deduced from the implied joint PDF $f_{X,Y}(x,y)$. From the same bivariate PDF, we can also deduce that $Y\mid X=x$ is gamma-distributed.
Proposition 2.
One can define a data-augmented Gibbs sampler for the GIG by sampling iteratively from the conditional distributions $X\mid Y=y$ and $Y\mid X=x$. Only the draws from $X\mid Y=y$ are relevant if we are interested in sampling from the GIG$(p,a,b)$.
In Propositions 1 and 2, we treat the cases $p<-1/2$ and $p>-1/2$ separately to obtain general results for all $p\neq-1/2$. For building a Gibbs sampler, one can focus on one of the cases and leverage the fact that if $X\sim\mathrm{GIG}(-p,b,a)$, then $1/X\sim\mathrm{GIG}(p,a,b)$ (Jørgensen, 1982).
The Gibbs sampler requires sampling from inverse Gaussian and gamma distributions. There are efficient algorithms to generate random draws from these distributions. For example, one can generate inverse Gaussian draws with the algorithm described in Michael et al. (1976), whereas for the gamma distribution, one can use standard algorithms such as those proposed in Ahrens and Dieter (1974).
The Gibbs sampler is geometrically ergodic. Geometric ergodicity is a property that quantifies the rate of convergence of the Gibbs sampler to its stationary distribution (see, for example, Johnson and Burbank (2015) for further details). It also ensures that inferences made with draws from the chain are well-behaved. In particular, if $\{X_{i}\}_{i=1}^{n}$ are draws from the chain and we want to estimate a function $f(X)$ with finite $\mathbb{E}[f(X)^{2+\varepsilon}]$ for some $\varepsilon>0$, geometric ergodicity ensures that $\sum_{i=1}^{n}f(X_{i})/n$ is approximately normal (Chan and Geyer, 1994). Our proof of geometric ergodicity relies on Theorem 3.5 in Johnson and Burbank (2015).
Proposition 3.
The data-augmented composition Gibbs sampler implied by Propositions 1 and 2 is geometrically ergodic.
The GIG$(p,a,b)$ often arises as a full conditional distribution in Gibbs sampling algorithms. There are countless examples from the literature on global-local shrinkage priors (Armagan et al., 2011), graphical models (Khare et al., 2018), stochastic volatility models (Barndorff-Nielsen, 1997), Bayesian nonparametrics (Favaro et al., 2012), smoothing splines (Loyal, 2024), and other areas. The Gibbs sampler can be embedded within an encompassing MCMC algorithm when simulation from a GIG$(p,a,b)$ is required. The following example offers a stylized illustration.
Example: Suppose that we observe independent and identically distributed data $y_{i}\mid\mu,\sigma^{2}\sim\mathrm{Normal}\left(\mu,\sigma^{2}\right)$, $i\in\{1,2,\,...\,,n\}$, with $\mu$ and $\sigma^{2}$ unknown. We assign independent priors to $\mu$ and $\sigma^{2}$ with $\mu\sim\mathrm{Normal}\left(\theta_{0},\tau_{0}^{2}\right)$ and $\sigma^{2}\sim\mathrm{GIG}\left(p_{0},a_{0},b_{0}\right).$ Letting $\mathbf{y}=\left(y_{1},\ldots,y_{n}\right)^{\top}$ and $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i},$ we can simulate from the posterior distribution of $\left(\mu,\sigma^{2}\right)$ by iterating between the full conditional distributions
where
and
If $p_{n}>-1/2,$ the full conditional distributions are
We conducted a simulation experiment to compare the efficiency of the two Gibbs samplers described in the example. We found that the data-augmented Gibbs sampler was more than 9 times faster than the original Gibbs sampler when the latter was implemented with the method of Devroye (2014) from the R package boodist (Laurent, 2023). However, the data-augmented Gibbs sampler was approximately 8% slower when the original Gibbs sampler was implemented with the method of Hörmann and Leydold (2014) from the R package GIGrvg (Leydold and Hormann, 2023). Neither the original Gibbs sampler nor the data-augmented Gibbs sampler produce independent draws from the posterior distribution, so it is important to evaluate the effective sample size (ESS) in addition to the time it takes to produce the Markov chains. For the original Gibbs sampler, we found that the ESS was close to the nominal sample size for both parameters $\mu$ and $\sigma^{2}.$ For the data-augmented Gibbs sampler, this was only true for $\mu.$ The ESS of $\sigma^{2}$ was about 1/3 of the nominal sample size. The details of the simulation experiment can be found in the supplementary material.
3 Exact sampling and CDF evaluation for half-integer $p$
The main results of this section reveal a relationship between GIGs with different values of $p.$ Proposition 4 leads to the first algorithm in the literature for exact, rejection-free simulation from the GIG$(p,a,b)$ for half-integer $p.$ Proposition 5 leads to an algorithm for evaluating the CDF of the GIG$(p,a,b)$ for half-integer $p.$ Pseudocode for the algorithms can be found in the supplementary material.
Proposition 4.
$f_{Y}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w\,f_{\mathrm{GIG}}(y\,;\,p-2,a,b)+(1-w)\,f_{\mathrm{GIG}}(y\,;\,p-1,a,b)}$
$w$ $={K_{p-2}(\sqrt{ab})}/{K_{p}(\sqrt{ab})}\in(0,1)$
and $E\sim\mathrm{Exp}(a/2)$, where $f_{\mathrm{GIG}}(y\,;\,p,a,b)$ is the PDF of the GIG$(p,a,b)$.
The proposition above can be proven with Theorem 2.4 of Jauch et al. (2023). It shows that if $p>1$ is half-integer, one can simulate from the $\mathrm{GIG}(p,a,b)$ by recursively simulating GIGs of smaller order $p$ until reaching the base case $p=3/2$. In that case, the components of the mixture are $\mathrm{GIG}(-1/2,a,b)$, which is equivalent to $\mathrm{InvGauss}(\sqrt{b/a},b)$. The case $\mathrm{GIG}(1/2,a,b)$ can be handled with an inverse Gaussian as well because if $X\sim\mathrm{GIG}(-p,b,a)$, then $1/X\sim\mathrm{GIG}(p,a,b)$. This property can also be used to include the cases where $p$ is a negative half integer.
The exact simulation method described above should be very efficient when $|p|$ is sufficiently small. To verify this, we conducted an experiment comparing the exact simulation method with the method Devroye (2014) from the R package boodist (Laurent, 2023) and the method of Hörmann and Leydold (2014) from the R package GIGrvg (Leydold and Hormann, 2023) in the case where $p=3/2.$ As expected, the exact simulation method outperformed the other two methods in terms of run time. The details of the simulation experiment can be found in the supplementary material.
Proposition 5.
where
and $\Gamma(\cdot,\cdot)$ is the incomplete gamma function (Abramowitz and Stegun, 1968).
The recurrence formula for $G_{p}$ implies a recursive algorithm for evaluating the CDF of the GIG$(p,a,b)$ for half-integer $p$. The base case involves the CDFs $G_{-1/2}$ and $G_{1/2}$, which are based on inverse Gaussian distributions.
It is not uncommon that statistical methods require sampling from or evaluating the CDF of the $\text{GIG}(p,a,b)$ for half-integer $p.$ Favaro et al. (2012) introduce a stick-breaking representation for the normalized inverse Gaussian process (Lijoi et al., 2005) whose practical value depends upon being able to efficiently simulate from the $\text{GIG}(p,a,b)$ for half-integer $p.$ In both Bhattacharya et al. (2015) and Loyal (2024), GIG$(p,a,b)$ full conditional distributions have half-integer $p$ parameters for the priors considered. He et al. (2022) present a data augmentation scheme for models that include gamma functions that requires simulating from a sum of independent GIGs for which $p=-3/2.$ Willmot and Woo (2022) describe several financial and actuarial applications in which one needs to evaluate the CDF of the $\text{GIG}(p,a,b)$ for half-integer $p.$
4 Conclusions
This article introduced two mixture representations of the GIG: one as a continuous mixture of inverse Gaussian distributions, and the other as a mixture of two GIGs plus an exponential random variable. These mathematical results led to conceptually straightforward algorithms for Monte Carlo simulation and CDF evaluation. As future work, it would be interesting to see if there are analogous representations for the matrix-variate version of the generalized inverse Gaussian (Barndorff-Nielsen et al., 1982; Hamura et al., 2024), a probability distribution over symmetric positive-definite matrices that includes the Wishart and inverse Wishart as special cases.
Appendix A Appendix
Section 1 includes the proofs for the propositions stated in the main text. Section 2 contains pseudocode for the algorithms. Section 3 considers an alternative parametrization of the GIG. Section 4 includes a derivation of a Gibbs sampler that can be easily deduced from results in Zhang and Reiter (2022). Finally, Section 5 includes a discussion on results from numerical experiments.
A.1 Proofs of Propositions
Proposition 1.
$f_{X}(x)=\int_{0}^{\infty}f_{X\mid Y}(x\mid y)f_{Y}(y)\,\mathrm{d}y,$
where $f_{X}(x)$ is the PDF of the $\mathrm{GIG}(p,a,b)$, $f_{Y}(y)$ is a PDF with support on $\mathbb{R}_{+}$, and $f_{X\mid Y}(x\mid y)$ is an inverse Gaussian PDF.
The parameters of the distributions are different for the cases $p<-1/2$ and $p>-1/2$. If $p<-1/2$, then
$f_{X\mid Y}(x)$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,=f_{\mathrm{IG}}(x\,;\,\sqrt{b/(a+2y)},b)}$
$\mathcal{Z}_{0}$ $=\frac{b^{(p+1)/2}\,\sqrt{2}\,\Gamma(-p-1/2)\,K_{p}(\sqrt{ab})}{a^{p/2}\,\sqrt{\pi}},$
where $f_{\mathrm{IG}}(x\,;\,\sqrt{b/(a+2y)},b)$ is the PDF of the InvGauss$(\sqrt{b/(a+2y)},b)$. If $p>-1/2$, then
$f_{X\mid Y}(x)$ $\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=f_{\mathrm{IG}}(x\,;\,\sqrt{(b+2y)/a},b+2y)}$
Proof.
$f_{X}(x)\propto x^{p+1/2}\,f_{\mathrm{IG}}(x\,;\,\sqrt{b/a},b),$
where $f_{\mathrm{IG}}(x\,;\,\mu,\lambda)$ is the PDF of an inverse Gaussian:
[…]
for $\mu>0$ and $\lambda>0$.
The term $x^{p+1/2}$ can be represented as follows
so we can write
$f_{X}(x)\propto f_{\mathrm{IG}}(x\,;\,\sqrt{b/a},b)\,\int_{0}^{\infty}y^{-(p+3/2)}\,\exp\left(-xy\right)\,\mathrm{d}y.$
If we define
$g(x,y)=f_{\mathrm{IG}}(x\,;\,\sqrt{b/a},b)\,y^{-(p+3/2)}\,\exp\left(-xy\right)\mathbbm{1}(x>0,y>0),$
$X\mid Y=y$ $\sim\mathrm{InvGauss}(\sqrt{b/(a+2y)},b)$
$Y\mid X=x$ $\sim\mathrm{Gamma}(-(p+1/2),x).$
$x^{p+1/2}\underset{(p>-1/2)}{=}\frac{\int_{0}^{\infty}y^{p-1/2}\,\exp(-y/x)\,\mathrm{d}y}{\Gamma(p+1/2)}.$
We can rewrite the PDF of the $\mathrm{GIG}(p,a,b)$ as
$f_{X}(x)\propto f_{\mathrm{IG}}(x\,;\,\sqrt{b/a},b)\,\int_{0}^{\infty}y^{p-1/2}\,\exp\left(-y/x\right)\,\mathrm{d}y.$
By Tonelli’s theorem, the joint PDF of $(X,Y)$ is proportional to
$f_{X,Y}(x,y)\propto y^{p-1/2}\exp\left\{-\sqrt{a(b+2y)}\right\}\,f_{\mathrm{IG}}(x\,;\,\sqrt{(b+2y)/a},b+2y).$
The conditional distributions are
$X\mid Y=y$ $\sim\mathrm{InvGauss}(\sqrt{(b+2y)/a},b+2y)$
$Y\mid X=x$ $\sim\mathrm{Gamma}(p+1/2,1/x).$
Lastly, we find $f_{Y}(y)$. Again, we treat the cases $p<-1/2$ and $p>-1/2$ separately. Let $f_{\mathrm{G}}(x\,;\,\alpha,\beta)$ be the PDF of the gamma distribution:
[…]
for $\alpha>0$ and $y>0$. If $p<-1/2$, then:
$f_{Y}(y)^{-1}$ $=\int_{0}^{\infty}\frac{f_{X}(x\mid y)}{f_{Y}(y\mid x)}\,\mathrm{d}x$
$=\int_{0}^{\infty}\frac{f_{\mathrm{IG}}(x\,;\,\sqrt{b/(a+2y)},b)}{f_{\mathrm{G}}(x\,;\,-(p+1/2),x)}\,\mathrm{d}x.$
Before we proceed with the integral, we write the explicit definitions of the densities. The inverse Gaussian density is
$f_{\mathrm{IG}}(x\,;\,\sqrt{b/(a+2y)},b)$ $=\frac{b^{1/2}x^{-3/2}}{(2\pi)^{1/2}}\,\exp\left\{\frac{-(a+2y)}{2x}\left[x^{2}-2x\sqrt{b/(a+2y)}+b/(a+2y)\right]\right\}$
$=\frac{b^{1/2}x^{-3/2}}{(2\pi)^{1/2}}\exp\left\{\,\frac{-(a+2y)x}{2}-\frac{b}{2x}+\sqrt{(a+2y)b}\right\}.$
The gamma density is
$f_{G}(y\,;\,-(p+1/2),x)=\frac{x^{-p-1/2}}{\Gamma(-p-1/2)}y^{-p-3/2}\exp\{-yx\}.$
Therefore,
$f_{Y}(y)^{-1}$ $=\int_{0}^{\infty}\frac{f_{\mathrm{IG}}(x\,;\,\sqrt{b/(a+2y)},b)}{f_{\mathrm{G}}(x\,;\,-(p+1/2),x)}\,\mathrm{d}x$
$=\mathcal{K}_{0}\int_{0}^{\infty}\underbrace{x^{p-1}\exp\left\{\frac{-(ax+b/x)}{2}\right\}}_{\propto\mathrm{GIG}(p,a,b)}\,\mathrm{d}x$
$=\mathcal{K}_{0}\frac{2K_{p}(\sqrt{ab})}{(a/b)^{p/2}}$
where
$\mathcal{K}_{0}=\frac{b^{1/2}\Gamma(-p-1/2)y^{p+3/2}\exp\{\sqrt{b(a+2y)}\}}{(2\pi)^{1/2}}.$
Rearranging normalizing constants, we obtain